OpenStudy (anonymous):

n/5 - 3n/10= 1/5 QB NEEDS HELP

OpenStudy (anonymous):

n=-2

OpenStudy (anonymous):

thankyou dany612!

OpenStudy (anonymous):

ok first find the common denominator, you do that by multiplying 2 to (n/5) then you get 2n/10-3n/10=-1n/10 set -1n/10=1/5 cross multiply -5n=10 divide by -5 n=-2

OpenStudy (anonymous):

you're welcome! and also thank amk11 for the step by step

OpenStudy (anonymous):

thank you amk11!:)

OpenStudy (anonymous):

no problem :)